OpenStudy (maheshmeghwal9):

Prove that the following function is continuous at {x=n}, where 'n' is a (+)ve integer :--

OpenStudy (maheshmeghwal9):

\[f(x)=x^n\]

OpenStudy (unklerhaukus):

well its derivatives are continuous

OpenStudy (unklerhaukus):

i dont know if that helps

OpenStudy (maheshmeghwal9):

What does it mean?

OpenStudy (maheshmeghwal9):

I mean derivatives of\[f(x)=x^n\]

OpenStudy (unklerhaukus):

\[f'(x)=nx^{n-1}\]

OpenStudy (maheshmeghwal9):

k!then, but how can we say that it is continuous?

OpenStudy (anonymous):

If it's not defined at n then it's it's not continuous.

OpenStudy (maheshmeghwal9):

But it is continuous 100%

OpenStudy (anonymous):

Did you prove it already?

OpenStudy (maheshmeghwal9):

no I didn't proved it ,but this is the question's demand!

OpenStudy (anonymous):

Do you get an exact value fo any positive n?

OpenStudy (anonymous):

Regardless of what positive you input you will always get an output.

OpenStudy (maheshmeghwal9):

no I got only \[f(n)=n^n\] where 'n' is a (+) integer

OpenStudy (anonymous):

continuos

OpenStudy (maheshmeghwal9):

& question's demand is to prove it by limit method

OpenStudy (maheshmeghwal9):

we have to do just like\[\lim_{x \rightarrow n^-} f(x)=\lim_{x \rightarrow n^+} f(x)=f(n)\] But I don't know how!

OpenStudy (anonymous):

continuous because \[\lim_{x\to a}x=a\] by sheer obviousness and therefore \[\lim_{x\to a}x^n=a^n\] because the limit of a product is the product of the limits

OpenStudy (maheshmeghwal9):

Then can we say that \[a^n\] where a is any real no., is continuous.If yes then why? please tell.

OpenStudy (anonymous):

as i understand it you have the function \(f(x)=x^n\) for some positive integer \(n\) and you want to show it is continuous (for any number \(a\)) this means you want to show that \[\lim_{x\to a}f(x)=f(a)\] in this case it is \[\lim_{x\to a}x^n=a^n\] that is what you need to show

OpenStudy (maheshmeghwal9):

k! now, can we show that f(x) is continuous at x=n from the above statement?

OpenStudy (anonymous):

and i wrote above what ought to convince anyone that it is true, unless you need to use a "epsilon, delta" proof

OpenStudy (anonymous):

well if it is true for all \(a\in \mathbb{R}\) is is certainly true for any \(n\)

OpenStudy (maheshmeghwal9):

k! thanx a lot.I got it at last.